OpenStudy (madison5s):

Find the axis of symmetry of the graph of the function. f(x) = 2x2 - 12x + 19

OpenStudy (madison5s):

@Will.H

OpenStudy (will.h):

Axis of symmetry = -b/2a Where b is -12 and a 2 Axis of symmetry = 12/4 Axis of symmetry = 3

OpenStudy (madison5s):

@Will.H THANK YOU SO MUCH!!!!!!!!!

OpenStudy (will.h):

You welcome :)